### v4

# The myth of declining competitive balance in the UEFA Champions League group stage

Abstract
The degree of competitive balance is a crucial issue in the analysis of sports competitions. According to previous studies, competitive balance has significantly declined in the UEFA Champions League group stage over the recent decades. Our paper introduces six alternative indices in order to explore the robustness of these results. The ex ante measures are based on Elo ratings, while our ex post measures compare the group ranking to reasonable benchmarks. We find no evidence of any trend in the competitive balance of the UEFA Champions League group stage between the 2003/04 and 2023/24 seasons.
“Different measurements are of different use, and all lines of research into competitive balance have, to date, proven quite instructive. To ignore this is to forgo important insights into the behavior of competitive balance.”111 Source: Fort and Maxcy, (2003, Abstract).
Keywords: competitive balance; Elo rating; football; Kendall rank correlation; UEFA Champions League
MSC class: 62P20, 91B82
JEL classification number: L11, Z20, Z21
1 Introduction
One of the most prestigious football tournaments, the UEFA Champions League, has been organised in the same format over 21 years, which has fundamentally been changed from the 2024/25 season. Commenting on this decision, the president of the Union of European Football Associations (UEFA), Aleksander Čeferin, said (UEFA,, 2022): “We are convinced that the format chosen strikes the right balance and that it will improve the competitive balance and generate solid revenues that can be distributed to clubs, leagues and into grassroots football across our continent while increasing the appeal and popularity of our club competitions.” Indeed, Gyimesi, (2024) has demonstrated that the new format can be successful in improving competitive balance.
Recent studies have observed a robust decline in both the ex ante (Triguero-Ruiz and Avila-Cano,, 2023) and ex post (Ramchandani et al.,, 2023; Triguero-Ruiz and Avila-Cano,, 2023) competitive balance of the UEFA Champions League group stage over the last decades, which might explain the reform. Our paper aims to check the sensitivity of these results by introducing and explaining alternative indicators of competitive balance. Naturally, the proposed measures may have their own disadvantages and shortcomings, and it is not claimed that they should be preferred to existing indices. However, it would be unreasonable to dismiss them for the sake of the standard measures of competitive balance since “there are no easy formulas for competitive balance success” (Zimbalist,, 2002, p. 120). We think it is the responsibility of the academic community to confirm or refute previous results on competitive balance by considering as many measures as possible, which will be done in the following.
In contrast to the conclusions of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023), no evidence is found of any decline in the competitive balance of the UEFA Champions League group stage in the recent decades, between the 2003/04 and 2023/24 seasons. Consequently, if UEFA has chosen the barely used incomplete round-robin design of the UEFA Champions League with its inherent risks (Devriesere et al.,, 2024) because of the worsening trend in competitive balance, the decision-makers might have been misled.
The paper has the following structure. Section 2 provides a concise literature review. Section 3 summarises some possible limitations of previous studies. Our measures of competitive balance are introduced in Section 4 together with a description of the underlying data. Section 5 provides the results and Section 6 discusses them. Finally, Section 7 offers concluding remarks.
2 Related literature
The economic analysis of team sports is often centred around the issue of competitive balance since the seminal articles of Rottenberg, (1956) and El-Hodiri and Quirk, (1971). Fort and Maxcy, (2003) distinguish two approaches. Analysis of competitive balance focuses on what has happened to competitive balance over time or as a result of rule changes. The second line of literature on competitive balance studies its effect on the stakeholders, especially fans, and usually involves testing the longstanding uncertainty of outcome hypothesis.
In recent decades, various indicators have been developed to quantify the competitive balance of sporting contests. The survey of Pawlowski and Nalbantis, (2019) classifies these measures according to the dimension of time. Short-term competitive balance refers to the uncertainty of individual games, often measured by using winning probabilities. The mid-term dimension focuses on seasonal uncertainty, while the long-term dimension concentrates on inter-seasonal uncertainty, the trend of competitive balance across several seasons. Our paper deals only with long-term competitive balance.
One commonly used measure of long-term competitive balance is the dispersion of winning percentage within sports leagues, also known as the Noll–Scully ratio (Noll,, 1988; Scully,, 1989). The Competitive Balance Ratio proposed by Humphreys, (2002) is an alternative measure that captures season-to-season changes in relative standings.
The Noll–Scully ratio is biased if leagues with different season lengths are compared. To address this issue, Doria and Nalebuff, (2021) have introduced new indices that are invariant to season length, which has led to changes in competitive balance rankings among some major sports leagues, with surprising results such as the NFL moving from the most balanced to the least balanced league. The current study offers a similar re-evaluation of a “common belief” on competitive balance in the UEFA Champions League.
The other popular concept to measure long-term competitive balance is the Herfindahl–Hirschman index () from the literature of industrial organization (Herfindahl,, 1950; Hirschman,, 1945, 1964), firstly used in sports by Depken, (1999). The normalised version of is even more widespread (Owen et al.,, 2007). Triguero Ruiz and Avila-Cano, (2019) construct a novel index based on the concept of distance, with a range in the unit interval, which is the square root of the normalised . It solves the problem of limited cardinality of most indices and makes the interpretation of the differences between the levels of competitive balance possible. Avila-Cano et al., (2021) determine the maximum concentration as a function of the number of teams and the scoring system. Analogously, Triguero-Ruiz et al., (2023) compute the minimal in sports leagues without ties.
Alternative indicators of long-term competitive balance have also been suggested. Groot, (2008, Chapter 1.4) proposes Kendall rank correlation to obtain a dynamic measure of competitive balance. Manasis et al., (2011) consider the normalised concentration ratio to evaluate the extent to which a league is dominated by a small subset of teams. Manasis et al., (2013) introduce an index to capture the degree of competition for winning the important prizes awarded in European football leagues such as the championship, the qualification for prestigious international club competitions, and avoiding relegation. Manasis and Ntzoufras, (2014) examine the applicability of between-season competitive balance indices in the context of European football, and define a set of specially designed measures. Hood and Jewell, (2022) illustrate the value of using betting data to simulate an ex-ante distribution of league-point outcomes in English football. Basini et al., (2023) develop a statistical network model to assess the balance between teams in a league. Csató and Petróczy, (2024) adopt bibliometric indices for quantifying competitive balance in the UEFA Champions League knockout stage. A recent empirical analysis of nontransitive patterns in professional football (van Ours,, 2024) can provide a novel approach to measuring competitive balance, too.
Traditional measures of performance have recently been challenged in multi-stage cycling races (Ausloos, 2024a, ; Ausloos, 2024b, ). Our paper has a similar aim in quantifying competitive balance in the group stage of (football) tournaments.
3 Potential limitations of previous studies
Competitive balance has two interpretations: ex ante and ex post. Ex ante competitive balance is experienced before the matches are played and is connected to suspense; the consumers hope to see an exciting game where winning and losing are far from being predetermined (Richardson et al.,, 2023). On the other hand, ex post competitive balance is related to surprise. According to Ely et al., (2015), it is more surprising if the current events are greatly different from the previous events.
Triguero-Ruiz and Avila-Cano, (2023) start to quantify ex ante competitive balance by calculating the shares of the clubs from UEFA club coefficients. They are used to compute for each group, which gives the distance to competitive balance () index (Dubois,, 2022; Scelles et al.,, 2022; Triguero Ruiz and Avila-Cano,, 2019) as follows:
UEFA club coefficient is the official measure of team strength in UEFA club competitions, the number of points gained by a team in UEFA club competitions over the last five years (Dagaev and Rudyak,, 2019). There is a lower bound based on the five-season association coefficient, which might be relevant for emerging clubs without meaningful experience at the international level (Csató,, 2022). Primarily, the rankings derived from the UEFA club coefficients have determined seeding in the UEFA club competitions (Csató,, 2020).
However, UEFA club coefficients suffer from a major shortcoming: they do not reflect the results of the majority of matches played by the clubs since all games in the national leagues and cups are omitted. Home advantage does not count either. Therefore, UEFA club coefficients usually underestimate emerging clubs in the top leagues without a recent international presence. A good example is the English club Aston Villa in the 2024/25 UEFA Champions League, which had the 31st highest UEFA club coefficient but was ranked 20th according to its Elo rating on 2 September 2024. In contrast, the Ukrainian champion Shakhtar Donetsk was 17th according to its UEFA club coefficient and 33rd according to its Elo rating. Unsurprisingly, a variant of the Elo method provides much higher accuracy in terms of explanatory power for the UEFA Champions League (Csató,, 2024).
Ex post competitive balance is usually quantified by the Herfindahl–Hirschman index () based on the shares of each team from the total points awarded (Avila-Cano et al.,, 2023; Owen and Owen,, 2022; Owen et al.,, 2007). Both Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023) use this approach to compute ex post competitive balance in a group. The underlying assumption is that the teams have appropriate incentives to collect points even though their qualification depends on the group ranking rather than on the number of points. Therefore, the clubs do not necessarily have appropriate incentives in the last round(s) (Chater et al.,, 2021; Csató et al.,, 2024). Buraimo et al., (2022) find robust evidence that matches of no importance for seasonal outcomes—which is equivalent to group ranking in our setting—may not be important to viewers in the English Premier League.
In the history of football, numerous examples exist when a team was satisfied with a (moderate) loss if it was still sufficient to achieve its objectives (Guyon,, 2020, 2022; Kendall and Lenten,, 2017). This consideration could be especially important in the UEFA Champions League group stage since the clubs usually play in their domestic leagues both on the previous and subsequent weekends, which creates a powerful incentive to rest the best players and field a younger squad to gain experience if the game is stakeless with respect to the rank of the team.
As an illustration, consider Group G in the 2018/19 UEFA Champions League (Csató et al.,, 2024). Before the last matchday, Real Madrid already won the group, and Roma was guaranteed to finish in the second place. In the last round, Real Madrid suffered a shocking 3-0 defeat against CSKA Moscow after fielding a fully rotated squad, which was its first European home loss by more than two goals (Bell,, 2018). Roma also lost against Viktoria Plzeň despite its home win of 5-0 in the second round. We have serious doubts that most consumers perceive the group to be more balanced—as implied by —merely because Real Madrid lost rather than won its last stakeless match.
Based on the arguments above, two indices of ex ante and four indices of ex post competitive balance will be constructed in order to check the robustness of previous findings by analysing alternative but reasonable indicators of competitive balance. Naturally, our measures have their own weaknesses, and they should not necessarily be preferred to the indicators used in the previous literature.
4 Methodology and data
The UEFA Champions League was organised in the same format between the 2003/04 and 2023/24 seasons. The group stage contained eight groups of four teams each. The top two teams qualified for the Round of 16, where the group winners were matched with the runners-up, the third-placed teams were transferred to the second most prestigious UEFA club competition, and the fourth-placed teams were eliminated.
In order to ensure the balancedness of the groups, the 32 teams were allocated into four pots. Until the 2014/15 season, seeding was determined by the UEFA club coefficients except for assigning the titleholder to the first pot (Csató,, 2021). The first pot consisted of the titleholder and the champions of the highest-ranked associations from the 2015/16 season (Dagaev and Rudyak,, 2019), and the UEFA Champions League titleholder, the UEFA Europa League titleholder, as well as the champions of the highest-ranked associations from the 2018/19 season (Csató,, 2020).
Sections 4.1 and 4.2 motivate and define the indices of ex ante and ex post competitive balance, respectively. Section 4.3 provides an example of their calculation, Section 4.4 gives information about the underlying data, while Section 4.5 presents how the dynamics of competitive balance is analysed. Last but not least, Section 4.6 summarises the main steps of our research.
4.1 Measures of ex ante competitive balance
Following Bosker and Gürtler, (2024) and Csató, (2024), we use Football Club Elo Ratings as a measure of team strength (Football Club Elo Ratings,, 2024). It gives the expected probability that team wins against team as:
where and are the Elo ratings of the two teams, respectively. Obviously, .
To update the Elo ratings, the expected winning probability is compared to the actual result (1 for win, 0.5 for draw, 0 for loss):
Since , the sum of Elo ratings does not change after any game. Finally, home advantage is taken into account by increasing the Elo difference in (1). The home advantage parameter is not fixed in advance but dynamically updated to guarantee that it converges to the actual effect of home advantage (Football Club Elo Ratings,, 2024).
Groups of four teams imply six pairs of clubs. For each pair, the winning probability of the stronger team—a value between 0.5 and 1—is computed according to (1). The six values are added to get , which is normalised similarly to the idea behind the :
is our first index of ex ante competitive balance, for which a lower value is preferred since it indicates a higher (expected) uncertainty in the group matches.
If only two teams qualify from a group of four, a group is usually said to be a harsh “group of death” with three strong teams, independent of the strength of the fourth team (Guyon,, 2015; Laliena and López,, 2019). Thus, it makes sense to focus on the three strongest teams and derive the indicator from the three pairs implied. Our second index of ex ante competitive balance, normalised to lie between 0 and 1 is
Naturally, always holds due to the definitions. However, we will only investigate the trends in the evolution of one particular measure, and never compare to directly.
It is worth noting that some papers have used similar indicators to quantify ex ante competitive balance. Hood and Jewell, (2022) adopt betting odds to simulate an ex-ante distribution of league outcomes. Gyimesi, (2024) uses the numerical average of the absolute Elo difference of each match to measure the uncertainty of match outcomes in the new format of the UEFA Champions League.
To conclude, our approach for quantifying ex ante competitive balance has two advantages: (1) it has a straightforward interpretation as the average (normalised) winning probability of the stronger team in each group match; and (2) since Elo ratings imply a higher predictive power than UEFA club coefficients, consumers may show greater sensitivity to its fluctuation within the groups. On the other hand, the number of teams per national association is allocated according to UEFA association coefficients, which are based on UEFA club coefficients. Furthermore, the competing teams are determined by the results of the different national leagues. Finally, the Elo rating is somewhat persistent and may not reflect the real strength of the team; however, this caveat also applies to the UEFA club coefficient.
Following Triguero-Ruiz and Avila-Cano, (2023), the ex ante measures of competitive balance and are averaged for the eight groups in each season.
4.2 Measures of ex post competitive balance
We compute the Kendall rank correlation coefficient (Kendall,, 1938) between an ex ante ranking and the final group ranking. The use of Kendall rank correlation for competitive balance goes back at least to Groot, (2008), who applied it to evaluate dynamic competitive balance in football leagues. In a sense, ex post competitive balance in the UEFA Champions League groups is about mobility: an ex ante (before the season) ranking of the teams can be compared to the final (after the season) group ranking to determine the uncertainty of tournament outcome. Mobility is maximal if the final ranking is exactly the reverse of the ex ante ranking, corresponding to a perfect negative rank correlation. However, then one can perfectly predict the final ranking by reversing the initial order of the teams. Hence, the ideal situation is if the final ranking is random, that is, Kendall rank correlation equals zero (Groot,, 2008).
Denote the number of concordant and discordant pairs between the ex ante and ex post rankings by and , respectively, and the number of teams by . The Kendall rank correlation is
The ex post ranking is, naturally, the final group ranking. However, two plausible choices exist for the ex ante rankings:
The initial pot allocation since each pot contains one team from each pot;
The ranking implied by the Football Club Elo Ratings before the season.
In addition, the consumers may ignore if the group winner and the runner-up are swapped because both of them qualify for the Round of 16. Then the maximum of discordant pairs is , and the analogously modified rank correlation measure is
where is the adjusted number of discordant pairs such that a discordant pair in the first two positions is disregarded, while all other discordant pairs are retained.
| Measure | Ex ante ranking | Are group winners and runners-up distinguished? |
| Pot allocation | ✔ (the formula of applies) | |
| Pot allocation | ✗ (the formula of applies) | |
| Elo rating | ✔ (the formula of applies) | |
| Elo rating | ✗ (the formula of applies) |
4.3 An illustrative example
The four teams of Group C in the 2023/24 UEFA Champions League are Napoli (drawn from Pot 1; Elo rating: 1911), Real Madrid (Pot 2; 1917), Braga (Pot 3; 1677), and Union Berlin (Pot 4; 1757). The pairwise differences between their Elo ratings are 6, 234, 154, 240, 160, 80, respectively, thus
Consequently, the stronger team wins a group match with a probability of 69% in average.
Analogously, the differences between the three highest Elo ratings are 6, 154, 160, respectively, thus
Consequently, the stronger team wins a group match with a probability of 64% in average if we focus on the three favourites.
The corresponding normalised values are:
The final group ranking is Real Madrid, Napoli, Braga, Union Berlin. Compared to the pot allocation, one discordant pair (Napoli and Real Madrid) exists, hence . However, this discordant pair affects the first two places in the group ranking, hence . Compared to the ranking implied by Elo ratings, there is one discordant pair (Braga and Union Berlin), hence . This discordant pair affects not only the first two places in the group ranking, hence .
4.4 Data
We consider all seasons between 2003/04 and 2023/24 when the tournament format of the UEFA Champions League did not changed. The pot allocation and group rankings are readily available from several sources; we have used Wikipedia after cross-checking with the official UEFA site.
Football Club Elo Ratings on a given date can be downloaded from api.clubelo.com/YYYY-MM-DD. In contrast to UEFA club coefficients, the Elo rating of a team changes during the season. Thus, the Elo ratings on 1 September are used for each season because the group stage is usually played between September and December. The Elo ratings are rounded to the nearest integer. Then Group B in the 2020/21 UEFA Champions League contains two teams (Shakhtar Donetsk and Borussia Mönchengladbach) with the same value (1766), which should be broken for indices and . It is decided by the higher value before rounding, but the tie-breaking does not affect our qualitative results.
4.5 The evolution of competitive balance
The approach of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023) is followed directly. For all the six (two ex ante and four ex post) indicators of competitive balance, the linear trend is obtained by estimating the following regression model:
where is the th competitive balance measure and is the error term in season (). Every season is denoted by its first year when the group stage is played. Furthermore, is the intercept and is the coefficient of the season. Competitive balance declined/did not changed/improved if is significantly positive/statistically zero/significantly negative.
4.6 Overview
Our results can be replicated by the following procedure:
1.
Collecting the publicly available Football Club Elo Ratings, UEFA Champions League pot allocations and group rankings between the 2003/04 and 2023/24 seasons (see Section 4.4 for details);
- 2.
3.
Estimation of linear regressions to check the existence of any trend in the seasonal averages of , , and – (see Section 4.5 for details).
5 Results
Notes: The seasons are denoted by their first year when the group stage is played.
The thick dashed red line shows the linear trend. None of them are significant even at a 25% significance level, see Table 2.
Figure 1 shows the evolution and the linear trend of the six competitive balance indices. The means of ex post competitive balance measures are around 0.5, and even the smallest seasonal average is above 0.2. Therefore, both pot allocation and Elo ranking are better predictors of the final group ranking than a random permutation (Groot,, 2008) as expected. Group G in the 2019/20 UEFA Champions League is an interesting outlier, where the final ranking is the opposite of the ranking implied by the pot allocation.
UEFA Champions League group stage between the 2003/04 and 2023/24 seasons
| Measure | Period | Intercept | Coefficient | -value | |
| 2003/04–2023/24 | 3.5418 | 0.001960 | 0.061 | 0.2784 | |
| 2003/04–2023/24 | 2.5841 | 0.001447 | 0.050 | 0.3297 | |
| 2003/04–2023/24 | 5.0573 | 0.002760 | 0.014 | 0.6047 | |
| 2003/04–2023/24 | 7.5424 | 0.004026 | 0.029 | 0.4574 | |
| 2003/04–2023/24 | 2.9361 | 0.001732 | 0.006 | 0.7460 | |
| 2003/04–2023/24 | 1.5176 | 0.001039 | 0.002 | 0.8583 |
Figure 1 may suggest a small drop in competitive balance over the recent decades since the slope of the linear trend is always positive. However, none of them are significant according to Table 2. This is in stark contrast with Triguero-Ruiz and Avila-Cano, (2023), where the -values are below 0.02 for ex ante and approximately 0 for ex post competitive balance, respectively. Unsurprisingly, never exceeds 0.1, contrary to the results of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023).
UEFA Champions League group stage between the 2003/04 and 2023/24 seasons
| Measure | Period | Intercept | Coefficient | -value | |
| 2003/04–2014/15 | 22.6219 | 0.01146 | 0.50 | 0.01034 | |
| 2014/15–2023/24 | 24.3567 | 0.01186 | 0.75 | 0.00124 | |
| 2003/04–2014/15 | 15.7854 | 0.00802 | 0.31 | 0.05861 | |
| 2014/15–2023/24 | 12.6759 | 0.00611 | 0.48 | 0.02656 |
Nonetheless, Table 3 presents that some trends can be found in ex ante competitive balance. In particular, it declined until the 2014/15 season and improved between the 2014/15 and 2023/24 seasons. The pattern is less robust but still exists if only the three strongest teams are considered in all groups as index does.
6 Discussion
How can these divergent conclusions be explained? Although both Ramchandani et al., (2023) (28 seasons from 1992/93 to 2019/20) and Triguero-Ruiz and Avila-Cano, (2023) (19 seasons from 1999/2000 to 2017/2018) study different samples, this cannot be the main reason. According to Ramchandani et al., (2023), the reduction in ex post competitive balance is even more evident between the 2002/03 and 2019/20 seasons, where they report for the linear regression; unfortunately, the -value is not reported. Analogously, the five-year averages of the DCB index continuously increase—the competitive balance worsens—in both the ex ante and ex post settings (Triguero-Ruiz and Avila-Cano,, 2023). Nonetheless, Table 3 suggests that the improvement in the recent seasons not considered in previous studies is partially responsible for the insignificant trend found here.
Our ex ante competitive balance indicators are based on the more accurate Elo rating. Triguero-Ruiz and Avila-Cano, (2023) use the UEFA club coefficients, which have likely become more concentrated among the leading clubs without a parallel dominance in the Elo ratings. A potential reason might be that the best clubs in the top leagues—except England, where the set of competitors is wider—can somewhat relax in their domestic leagues now since the number of guaranteed slots in the UEFA Champions League group stage has increased from two to four for the four highest-ranked leagues. Therefore, they are able to increasingly focus on international competitions.
Regarding ex post competitive balance, we compare the final group ranking to a benchmark provided by either the pot allocation or the Elo ranking. Since the does not require such a benchmark, a worsening ex post competitive balance may be covered if our benchmark becomes a more accurate predictor over time. However, this is improbable. Even though the seeding system has been reformed as discussed in Section 3, it has rather increased uncertainty (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). Similarly, the differences in the Elo ratings have not increased (Figure 1), thus, the quality of Elo ranking seems to be unchanged. On the other hand, the clubs may focus in the group stage more strongly on their number of points than previously, possibly due to the increasing financial incentives provided by the UEFA. This leads to a higher and a lower probability of losing stakeless games. Then the earlier studies have only mixed a favourable trend (teams do not lose their stakeless matches) with declining ex post competitive balance. The hypothesis may be tested by an empirical investigation of stakeless games, which can be identified, for example, by the approach of Csató et al., (2024). Finally, it should be noted that the sample size is rather small. However, the novel format of the UEFA Champions League from the 2024/25 season does not allow for more observations.
7 Conclusions
UEFA has chosen a risky strategy by fundamentally reforming the well-established group stage of its major tournament, the UEFA Champions League, from the 2024/25 season. According to the existing literature based on straightforward -based measures, UEFA should have acted effectively because of the notable drop in both the ex ante and ex post competitive balance of the group stage over the last two decades. Our paper has proposed six alternative but reasonable indices to see the robustness of these results. Crucially, no evidence is found for a declining competitive balance between the 2003/04 and 2023/24 seasons. Therefore, we urge researchers and analysts to consider a broader set of competitive balance indicators in the future before proposing powerful policy interventions.
Acknowledgements
This paper could not have been written without Réka Boros and Adrienn Czakó, who prepared the data and made the calculations.
We are indebted to András Gyimesi for useful suggestions.
One anonymous reviewer provided useful comments on an earlier draft.
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838 and PD 146055, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.